Express V= 3t² + 7t + -4 as a linear combination of the polynomials

P1= t² + 2t + 3

P2= 2t² + 3t + 7

P3 = 3t² + 5t + 6


1
Expert's answer
2021-08-02T14:46:08-0400

Let us express "V= 3t^2+ 7t -4" as a linear combination of the polynomials

"P_1= t^2 + 2t + 3,\ 

P_2= 2t^2 + 3t + 7,\ 

P_3 = 3t^2 + 5t + 6."


Let "V=aP_1+bP_2+cP_3."

Then

"3t^2+ 7t -4= a(t^2 + 2t + 3)+b( 2t^2 + 3t + 7)+c( 3t^2 + 5t + 6)."

It follows that

"3t^2+ 7t -4= at^2 + 2at + 3a+2bt^2 + 3bt + 7b+3ct^2 + 5ct + 6c,"

and hence

"3t^2+ 7t -4= (a+2b+3c)t^2 + (2a+3b+5c)t + (3a+ 7b + 6c)."


Therefore, we have the following system:


"\begin{cases}
a+2b+3c=3\\ 2a+3b+5c=7\\ 3a+ 7b + 6c=-4
\end{cases}"


Let us multiply the first equation by "-2" and add to the second equation, also multiply the first equation by "-3" and add to the third equation. Thgen we have the system:


"\begin{cases}
a+2b+3c=3\\ -b-c=1\\ b -3c=-13
\end{cases}"


Let us add the last to equation. Then we get the system:


"\begin{cases}
a=3-2b-3c\\ -b-c=1\\ -4c=-12
\end{cases}"


The last system is equivalent to the system:


"\begin{cases}
a=2\\ b=-4\\ c=3
\end{cases}"


We conclude that "V=2P_1-4P_2+3P_3."